littlebear17:

A mountain climber hikes 25.5 km from her base camp, going 20.0º north of east. On the second day, she walks 41.0 km in a direction 65.0º north of east, at which point she has reached a summit. Determine the magnitude and direction of her resultant displacement between the base camp and the second camp. (Remember to analyze the two motions separately, and then add and use Pythagorean theorem for the resultant and trigonometry for the angle.)

Vocaloid:

Decompose each hike into its x and y components x components = 25.5cos(20) + 41cos(65) y components = 25.5cos(20) + 41sin(65) to get the magnitude calculate the x and y components then take Magnitude = sqrt(x^2+y^2) Direction = arctan(y/x) As the problem states these are both derived from the pythagorean theorem/ basic trig